# Real-Time Cell Sorting with Scalable In Situ FPGA-Accelerated Deep Learning

### v1

@@ Abstract
Precise cell classification is essential in biomedical diagnostics and therapeutic monitoring, particularly for identifying diverse cell types involved in various diseases. Traditional cell classification methods such as flow cytometry depend on molecular labeling which is often costly, time-intensive, and can alter cell integrity. To overcome these limitations, we present a label-free machine learning framework for cell classification, designed for real-time sorting applications using bright-field microscopy images. This approach leverages a teacher-student model architecture enhanced by knowledge distillation, achieving high efficiency and scalability across different cell types. Demonstrated through a use case of classifying lymphocyte subsets, our framework accurately classifies T4, T8, and B cell types with a dataset of 80,000 preprocessed images, accessible via an open-source Python package for easy adaptation. Our teacher model attained 98% accuracy in differentiating T4 cells from B cells and 93% accuracy in zero-shot classification between T8 and B cells. Remarkably, our student model operates with only 0.02% of the teacher model’s parameters, enabling field-programmable gate array (FPGA) deployment. Our FPGA-accelerated student model achieves an ultra-low inference latency of just 14.5 s and a complete cell detection-to-sorting trigger time of 24.7 s, delivering 12x and 40x improvements over the previous state-of-the-art real-time cell analysis algorithm in inference and total latency, respectively, while preserving accuracy comparable to the teacher model. This framework provides a scalable, cost-effective solution for lymphocyte classification, as well as a new SOTA real-time cell sorting implementation for rapid identification of subsets using in situ deep learning on off-the-shelf computing hardware.
Abstract
Precise cell classification is essential in biomedical diagnostics and therapeutic monitoring, particularly for identifying diverse cell types involved in various diseases. Traditional cell classification methods such as flow cytometry depend on molecular labeling which is often costly, time-intensive, and can alter cell integrity. To overcome these limitations, we present a label-free machine learning framework for cell classification, designed for real-time sorting applications using bright-field microscopy images. This approach leverages a teacher-student model architecture enhanced by knowledge distillation, achieving high efficiency and scalability across different cell types. Demonstrated through a use case of classifying lymphocyte subsets, our framework accurately classifies T4, T8, and B cell types with a dataset of 80,000 preprocessed images, accessible via an open-source Python package for easy adaptation. Our teacher model attained 98% accuracy in differentiating T4 cells from B cells and 93% accuracy in zero-shot classification between T8 and B cells. Remarkably, our student model operates with only 0.02% of the teacher model’s parameters, enabling field-programmable gate array (FPGA) deployment. Our FPGA-accelerated student model achieves an ultra-low inference latency of just 14.5 s and a complete cell detection-to-sorting trigger time of 24.7 s, delivering 12x and 40x improvements over the previous state-of-the-art (SOTA) real-time cell analysis algorithm in inference and total latency, respectively, while preserving accuracy comparable to the teacher model. This framework provides a scalable, cost-effective solution for lymphocyte classification, as well as a new SOTA real-time cell sorting implementation for rapid identification of subsets using in situ deep learning on off-the-shelf computing hardware.
1 Introduction
@@
Accurate cell classification is critical for a wide range of biomedical applications, including disease diagnostics, immunological studies, and personalized therapies. Traditional methods for cell classification, such as molecular labeling through flow cytometry, rely on detecting specific surface markers [1]. While these techniques are accurate, they have notable limitations, including high costs, time-intensive protocols, and potential interference with the natural state of the cells [2]. In response to these challenges, label-free classification methods have emerged as a promising alternative by leveraging intrinsic cell properties, such as morphology and biomechanics. These approaches preserve the natural state of the cells, enabling downstream applications such as transplantation, functional studies, and real-time analysis or sorting [3].
@@
Recent advancements in Machine Learning (ML) have revolutionized cell classification by offering innovative solutions to circumvent the limitations of traditional methods. Despite their potential, many ML models perform suboptimally on datasets other than those they were specifically trained on, revealing inadequate generalization and transfer learning capabilities. Furthermore, training protocols optimized for general image datasets often fail to translate effectively to biological datasets [4, 5].
@@
Addressing these challenges requires robust training methodologies tailored specifically for diverse biological image datasets. In this study, we focus on optimized training protocols that achieve high specificity and sensitivity in cell classification. Using lymphocyte classification as a use case, we demonstrate the adaptability and effectiveness of these training recipes, highlighting their potential to extend seamlessly to various cell types and enabling versatile applications across different biological contexts. Specialized expertise in lymphocyte classification remains limited even in well-resourced communities, leading to variability in diagnostic accuracy. This issue is exacerbated in underserved areas, where the lack of access to expert pathology services results in prolonged or erroneous diagnostic outcomes that critically impair patient management. Our ML framework leverages bright-field images to detect cellular morphological features for the cell classification process. By eliminating reliance on molecular labels, this approach reduces human subjectivity, ensures reproducibility, and offers consistent results across different settings. Both our trained models and the detailed training recipes are publicly available through GitHub [6].
@@
Moreover, to meet the demands of real-time inference, we have implemented a field-programmable gate array (FPGA) version of our optimized student model, achieving ultra-low latency and high throughput. Previous studies have demonstrated cell classification machine learning inference performance on the order of milliseconds, primarily on GPU and CPU hardware [7, 8, 9]. The previous state-of-the-art (SOTA) in terms of inference latency implements a deep neural network (DNN) for standing surface acoustic wave cell sorting and achieves an inference latency of approximately 183 s [2] and a full cell detection-to-sorting trigger latency of <1 ms. With the advent of new tools like High-Level Synthesis (HLS) for Machine Learning (hls4ml) for neural network inference on FPGAs, we institute a new SOTA benchmark for cell classification.
@@
By leveraging our ML framework in a use case involving the classification of T4, T8, and B cells, we have achieved remarkable accuracy improvements. Our teacher model demonstrates approximately 98% accuracy in classifying T4 cells from B cells and achieves about 93% accuracy in zero-shot classification of T8 vs. B cells. Employing Knowledge Distillation (KD) techniques, our student 2 model attains sufficiently high accuracy relative to the teacher model with just about 0.02% of its parameters. The FPGA implementation of the student model further enhances processing speed, reducing inference latency to just 14.5 s. This improvement in processing speed facilitates the real-time analysis and accurate sorting of T and B cells, significantly advancing their rapid classification in clinical settings. With these insights and results in place, the core achievements and contributions of our study are summarized in the following research highlights:
@@
Dataset: We present a dataset of 80,000 images, which supports the training and validation of our models. The data can be seamlessly downloaded following the installation of our pip-installable package[10].
@@
Models: We provide detailed recipes for tuning a larger teacher model to predict T4 versus B cell classifications and training a smaller student model through KD for a 15X reduction in the inference speed with 0.02% of the computational resources.
@@
Transfer Learning: We demonstrate the transfer learning capability for T8 versus B cell classifications, indicating that the model can perform zero-shot inference and can be further tuned to detect other lymphocyte cell types.
@@
In Situ FPGA Implementation: We deploy our student model in situ on frame grabber FPGA hardware to elminate data transfer overhead and reduce inference latency from the 183 s previous SOTA and 325 s on GPU to just 14.5 s, a 12X and 22X improvement, respectively. Thus we institute a new SOTA real-time deep learning benchmark and implementation for real-time cell sorting and rapid classification.
@@ 2.3 Detection of Cell Class by Student:
Our experiments revealed that data-mixing augmentation techniques, such as CutMix and MixUp, substantially enhance KD performance. Conversely, other image-based augmentations, including random flipping and shearing, degraded the accuracy of the distilled student model [20]. Maintaining identical image crops and augmentation strategies for both teacher and student networks during training was crucial to ensure consistent learning and effective knowledge transfer without misalignment in data representation [19].
@@ 2.5 FPGA Implementation of the Student Model:
We then empirically benchmark the latency of the FPGA implementation of Student 2 by monitoring the internal communication protocol used by the neural network intellectual property (IP). We then utilize the frame grabber’s TTL IO to output a square wave where the high time denotes inference latency which we measure with an oscilloscope. Figure 6a exhibits the results of this latency test, showing a model inference latency of just 14.5 s. Additionally, we observe that inference begins approximately 10.0 s after the trigger edge. Given a 2 s exposure time, our model completes inferencing approximately 22.5 s after image exposure is finished. The model output writeout procedure takes an additional 0.2 s. The writeout consists of the model’s two-bit output indicating the cell output class, and can be expanded or adapted for any cell classification task or communication protocol. Aggregating these constituent components yields a full cell detection-to-sorting trigger time of 24.7 s. As shown in Figure 6b, we pipeline neural network inference with the exposure and readout processes to accelerate the algorithm to a throughput of 81 kfps in our implementation. This benchmark far exceeds our GPU’s best performance at a batch size of 1. Note that in Figure 6a we capture at 50 kfps such that consecutive inference traces do not overlap for readability purposes.
@@ 5 Data Availability
The LymphoMNIST dataset described in this work is publicly available to promote reproducibility and facilitate further research. Researchers can readily access and integrate the dataset into their workflows using the Python package LymphoMNIST. Comprehensive documentation, installation guidelines, dataset exploration methods, and implementation examples are provided in the project’s GitHub repository (https://github.com/Khayrulbuet13/LymphoMNIST).
@@ 6 Code Availability
All models and experiments described in this work were implemented using Python with PyTorch as the core deep learning framework. Complete scripts, model training recipes, and instructions necessary to reproduce the experiments and results reported in this study are openly accessible via https://github.com/Khayrulbuet13/LymphoML.

### v2

@@ Abstract
Precise cell classification is essential in biomedical diagnostics and therapeutic monitoring, particularly for identifying diverse cell types involved in various diseases. Traditional cell classification methods, such as flow cytometry, depend on molecular labeling, which is often costly, time-intensive, and can alter cell integrity. Real-time microfluidic sorters also impose a sub-ms decision window that existing machine-learning pipelines cannot meet. To overcome these limitations, we present a label-free machine learning framework for cell classification, designed for real-time sorting applications using bright-field microscopy images. This approach leverages a teacher–student model architecture enhanced by knowledge distillation, achieving high efficiency and scalability across different cell types. Demonstrated through a use case of classifying lymphocyte subsets, our framework accurately classifies T4, T8, and B cell types with a dataset of 80,000 pre-processed images, released publicly as the LymphoMNIST package for reproducible benchmarking. Our teacher model attained 98% accuracy in differentiating T4 cells from B cells and 93% accuracy in zero-shot classification between T8 and B cells. Remarkably, our student model operates with only 5682 parameters (0.02% of the teacher, a 5000-fold reduction), enabling field-programmable gate array (FPGA) deployment. Implemented directly on the frame-grabber FPGA as the first demonstration of in-situ deep learning in this setting, the student model achieves an ultra-low inference latency of just 14.5 s and a complete cell detection-to-sorting trigger time of 24.7 s, delivering 12 and 40 improvements over the previous state of the art in inference and total latency, respectively, while preserving accuracy comparable to the teacher model. This framework establishes the first sub-25 s ML benchmark for label-free cytometry and provides an open, cost-effective blueprint for upgrading existing imaging sorters.
keywords:
CNN, Cell Sorting, FPGA, Computer Vision, Knowledge Distillation1 Introduction
@@
Accurate cell classification is critical for a wide range of biomedical applications, including disease diagnostics, immunological studies, and personalized therapies. Traditional methods for cell classification, such as molecular labeling through flow cytometry, rely on detecting specific surface markers [1]. While these techniques are accurate, they have notable limitations, including high costs, time-intensive protocols, and potential interference with the natural state of the cells [2]. Equally important, modern acoustofluidic sorters provide only a 1 ms window between image acquisition and actuation, a latency budget that no published machine-learning (ML) pipeline has yet satisfied[3]. In response to these challenges, label-free classification methods have emerged as a promising alternative by leveraging intrinsic cell properties, such as morphology and biomechanics. Recent work has demonstrated the fundamental interconnection between biophysical cues and cellular morphology, with substrate geometry alone capable of reverting pluripotent stem cells to naivety through morphological changes[4]. This highlights that morphological features captured in bright-field images encode meaningful information about cellular state and phenotype. These approaches preserve the natural state of the cells, enabling downstream applications such as transplantation, functional studies, and real-time analysis or sorting [5, 6].
@@
Recent advancements in ML have revolutionized cell classification by offering innovative solutions to circumvent the limitations of traditional methods. For instance, deep CNNs have been successfully applied to bright-field images for label-free identification of cell types, with feature fusion approaches integrating morphological patterns across multiple convolutional modules to achieve high accuracy [7]. While such specialized approaches show promise, many general ML models perform suboptimally on datasets other than those they were specifically trained on, revealing inadequate generalization and transfer-learning capabilities. Furthermore, training protocols optimized for general image datasets often fail to translate effectively to biological datasets [8, 9]. Progress is also slowed by the scarcity of large, publicly available bright-field datasets with ground-truth phenotypes, making reproducible benchmarking difficult.
@@
Addressing these challenges requires robust training methodologies tailored specifically for diverse biological image datasets. In this study, we focus on optimized training protocols that achieve high specificity and sensitivity in cell classification. Using lymphocyte classification as a use case, we demonstrate the adaptability and effectiveness of these training recipes, highlighting their potential to extend seamlessly to various cell types and enabling versatile applications across different biological contexts. Specialized expertise in lymphocyte classification remains limited even in well-resourced communities, leading to variability in diagnostic accuracy. This issue is exacerbated in underserved areas, where the lack of access to expert pathology services results in prolonged or erroneous diagnostic outcomes that critically impair patient management. Our ML framework leverages bright-field images to detect cellular morphological features for the cell classification process. By eliminating reliance on molecular labels, this approach reduces human subjectivity, ensures reproducibility, and offers consistent results across different settings. To facilitate community adoption, we release both our training code and the 80 000-image LymphoMNIST dataset as pip-installable packages.
@@
Moreover, to meet the demands of real-time inference, we have implemented a field-programmable gate array (FPGA) version of our optimized student model, achieving ultra-low latency and high throughput. Previous studies have demonstrated cell classification ML inference performance on the order of milliseconds, primarily on GPU and CPU hardware [10, 11, 12]. The previous state-of-the-art (SOTA) in terms of inference latency implements a deep neural network (DNN) for standing surface acoustic wave cell sorting and achieves an inference latency of approximately 183 µs [2] and a full cell detection-to-sorting trigger latency of ms. Leveraging high-level-synthesis tools (hls4ml) and a knowledge-distilled student network with only 5682 parameters (about 0.02 % of the 28 M-parameter teacher, a 5000-fold reduction), we achieve the first frame-grabber-resident deep-learning implementation that fits within this strict latency envelope.
@@
By leveraging our ML framework in a use case involving the classification of T4, T8, and B cells, we have achieved remarkable accuracy improvements. Our teacher model demonstrates approximately 98 % accuracy in classifying T4 cells from B cells and achieves about 93 % accuracy in zero-shot classification of T8 vs. B cells. Employing knowledge-distillation (KD) techniques, our Student 2 model attains sufficiently high accuracy relative to the teacher model with just about 0.02 % of its parameters. The FPGA implementation of the student model further enhances processing speed, reducing inference latency to just 14.5 µs. This improvement in processing speed facilitates the real-time analysis and accurate sorting of T and B cells, significantly advancing their rapid classification in clinical settings. With these insights and results in place, the core achievements and contributions of our study are summarized in the following research highlights:
@@
Dataset: We present a dataset of 80000 images, which supports the training and validation of our models. The data are freely available via the pip-installable LymphoMNIST package for immediate benchmarking.
@@
Models: We publish detailed recipes for a high-capacity teacher and a KD-trained student with an in-house, lightweight architecture tuned for bright-field cell images, achieving 5000-fold parameter compression (5682 params, 0.02 % of the teacher) while retaining F1 0.97.
@@
Transfer Learning: We demonstrate the transfer-learning capability for T8 versus B cell classifications, indicating that the model can perform zero-shot inference and can be further tuned to detect other lymphocyte cell types.
@@
In-Situ FPGA Implementation: We deploy our student model directly on the frame-grabber FPGA, eliminating PCIe transfer overhead and reducing inference latency from the 183 µs previous SOTA and 325 µs on GPU to just 14.5 µs, a 12× and 22× improvement, respectively. Thus, we institute a new SOTA real-time deep-learning benchmark and implementation for real-time cell sorting and rapid classification.
@@ 2.3 Detection of Cell Class by Student:
Our experiments reconfirmed that data-mixing augmentation techniques, such as CutMix and MixUp, substantially enhance KD performance. Conversely, other image-based augmentations, including random flipping and shearing, degraded the accuracy of the distilled student model when applied inconsistently between teacher and student [23], as demonstrated by Beyer et al. [22]. Maintaining identical image crops and augmentation strategies for both teacher and student networks during training was crucial to ensure consistent learning and effective knowledge transfer without misalignment in data representation [22].
@@ 2.5 FPGA Implementation of the Student Model:
We then empirically benchmark the latency of the FPGA implementation of Student 2 by monitoring the internal communication protocol used by the neural network intellectual property (IP). We then utilize the frame grabber’s TTL IO to output a square wave where the high time denotes inference latency which we measure with an oscilloscope. Figure 6a exhibits the results of this latency test, showing a model inference latency of just 14.5 s. Additionally, we observe that inference begins approximately 10.0 s after the trigger edge. Given a 2 s exposure time, our model completes inferencing approximately 22.5 s after image exposure is finished. The model output writeout procedure takes an additional 0.2 s. The writeout consists of the model’s two-bit output indicating the cell output class, and can be expanded or adapted for any cell classification task or communication protocol. Aggregating these constituent components yields a full cell detection-to-sorting trigger time of 24.7 s. By reducing inference latency to under 25 s, our pipeline shifts the limiting factor from computation to fluidics. This margin not only exceeds the 1 ms actuation window of current acoustofluidic sorters [3], but also opens the door to applications previously inaccessible to image-based ML—such as sorting extracellular vesicles or bacteria, where transit times are an order of magnitude shorter than for mammalian cells. As shown in Figure 6b, we pipeline neural network inference with the exposure and readout processes to accelerate the algorithm to a throughput of 81 kfps in our implementation. This benchmark far exceeds our GPU’s best performance at a batch size of 1. Note that in Figure 6a we capture at 50 kfps such that consecutive inference traces do not overlap for readability purposes.
@@ 5 Data Availability
The LymphoMNIST dataset described in this work is publicly available to promote reproducibility and facilitate further research [40]. Researchers can readily access and integrate the dataset into their workflows using the Python package LymphoMNIST. Comprehensive documentation, installation guidelines, dataset exploration methods, and implementation examples are provided in the project’s GitHub repository (https://github.com/Khayrulbuet13/LymphoMNIST).
@@ 6 Code Availability
All models and experiments described in this work were implemented using Python with PyTorch as the core deep learning framework. Complete scripts, model training recipes, and instructions necessary to reproduce the experiments and results reported in this study are openly accessible [41] via https://github.com/Khayrulbuet13/LymphoML.
@@ 7 Acknowledgements
S2 FPGA Resource-Latency Trade-offs and Scalability Limits
The reuse factor hyperparameter in hls4ml controls the degree of resource sharing in the synthesized hardware, directly affecting both latency and resource consumption. This trade-off can be visualized as a Pareto frontier, where design points represent different balance points between parallelization (low reuse factor, low latency, high resources) and resource efficiency (high reuse factor, high latency, low resources). Figure SS1 illustrates this trade-off for neural network FPGA implementations of comparable complexity to our Student 2 model. As reuse factor increases from 2 to 1024, latency increases from tens of microseconds to milliseconds while resource consumption (measured as percentage of available LUTs, registers, BRAMs, and DSPs on the KU035 FPGA) decreases dramatically. At reuse factor = 2, DSP utilization approaches 100%, representing the low-latency extreme of the feasible design space. Our Student 2 implementation operates at this frontier, using reuse factors of 1–2 for convolutional layers and 25 for dense layers, achieving 14.5 µs inference latency while consuming approximately 95% of available DSPs.
Several factors constrain scalability to larger models or multi-class extensions on the target KU035 device. First, DSP exhaustion is the primary bottleneck: our design already consumes 95% of available DSP blocks at near-minimum reuse factors. Extending to deeper networks or higher-resolution inputs (e.g., 6464 instead of 4848) would require either increasing reuse factors—thereby trading latency for feasibility—or adopting LUT-based arithmetic, which increases LUT consumption and may reduce maximum clock frequency. Second, memory bandwidth and capacity impose practical limits: aggressive parallelization demands more concurrent memory ports and greater BRAM partitioning, which can exceed on-chip resources for large models. Third, routing congestion becomes significant in highly parallelized designs, potentially degrading such that even reduced cycle counts do not proportionally improve end-to-end latency. Finally, multi-class extensions increase computational and memory demands: for example, a 10-class classifier would require approximately 5 more DSPs and parameters in the final layer (a consequence of adding 8 more neurons) compared to our binary design. Our Student 2 architecture and heterogeneous reuse factor strategy represent a carefully optimized design point that achieves ultra-low latency (14.5 µs inference, 24.7 µs total detection-to-trigger) while remaining within the strict resource envelope of the KU035 FPGA. This positions our implementation at the practical performance limit for this device class and demonstrates the necessity of co-designing model architecture and hardware configuration to meet real-time sorting constraints.
S3 Model Calibration Analysis
Probability calibration measures whether a model’s predicted confidence scores accurately reflect its true classification accuracy. A well-calibrated model assigns high probabilities to samples it classifies correctly and low probabilities to potential misclassifications. We evaluated calibration performance using Expected Calibration Error (ECE) and Maximum Calibration Error (MCE) with 5-bin reliability analysis [Guo2017-rk]. ECE measures the average difference between predicted confidence and actual accuracy across bins, while MCE captures the worst-case deviation. Temperature scaling [Guo2017-rk] was applied as a post-hoc calibration technique to assess whether simple rescaling of the output logits could improve calibration metrics.
Figure S2 presents reliability diagrams for the Teacher and Student 1 models before and after temperature scaling. The Teacher model demonstrates excellent baseline calibration (ECE = 0.90%, MCE = 9.58%), with post-hoc temperature scaling (optimal temperature ) further improving ECE to 0.41%—a 54.7% reduction. The Student 1 model exhibits moderate baseline calibration (ECE = 1.59%, MCE = 7.42%) and displays underconfidence across low-to-mid confidence bins, meaning it tends to assign lower probabilities than its actual accuracy warrants. This conservative behavior is advantageous for classification tasks where false positives carry higher costs than false negatives, as underconfident predictions naturally provide a safety margin. Temperature scaling reduces worst-case calibration error (MCE from 7.42% to 4.75%), but increases average calibration error (ECE to 2.51%), indicating that attempts to shift predictions toward perfect calibration disrupt the model’s inherent conservative bias.
S4 Uncertainty-Aware Rejection and Throughput-Safety Trade-offs
Real-world deployment of machine learning models for cell sorting requires careful consideration of performance under domain shift and the ability to reject ambiguous predictions. In clinical settings, false routing of cells to incorrect collection channels can compromise downstream analyses or therapies. To address this critical deployment concern, we implemented and evaluated an uncertainty-aware rejection framework that quantifies the trade-off between throughput (fraction of samples processed) and safety (accuracy on processed samples).
Our rejection mechanism is based on confidence thresholding, where samples with maximum softmax probability below a threshold are rejected and not processed (directed to a discard channel in a physical sorting implementation). We evaluated seven threshold levels: on both our Teacher and Student1 models. To simulate domain shift conditions that would be encountered when deploying the system across different microscopes, illumination conditions, or sample preparation protocols, we applied realistic augmentation transforms to our test dataset. These transforms included color jitter with brightness and contrast variations of 0.4 (simulating illumination changes), random rotations up to 30 degrees (simulating sample orientation variations), random horizontal flips with 50% probability (simulating microscope positioning differences), and Gaussian blur (simulating optical variations). The test dataset consisted of 7,140 samples from the LymphoMNIST dataset, using a 70% test split with random state 42 for reproducibility.
| Model | Threshold | Coverage | Clean Acc. | Shift Acc. | False-Route Rate |
| (%) | (%) | (%) | (Clean / Shift, %) | ||
| Teacher | 0.50 | 100.0 | 97.4 | 94.6 | 2.6 / 5.4 |
| 0.85 | 94.6 | 98.9 | 97.9 | 1.1 / 2.1 | |
| 0.95 | 89.4 | 99.5 | 99.0 | 0.5 / 1.0 | |
| Student1 | 0.50 | 100.0 | 94.5 | 84.4 | 5.5 / 15.6 |
| 0.85 | 90.2 | 97.6 | 94.6 | 2.4 / 5.4 | |
| 0.95 | 83.3 | 98.8 | 97.4 | 1.2 / 2.6 |
Table S2 presents the comprehensive results of our uncertainty-aware rejection analysis. At the baseline threshold (), both models process 100% of samples. The Teacher model achieves 97.4% accuracy on clean data and 94.6% under domain shift, corresponding to a degradation of 2.8%. Student1 achieves 94.5% accuracy on clean data and 84.4% under shift, showing a 10.1% degradation. This baseline comparison confirms that the Teacher model exhibits superior robustness to distribution shift.
As we increase the confidence threshold, both models demonstrate the expected trade-off: coverage decreases while accuracy on processed samples increases. At the intermediate threshold (), the Teacher processes 94.6% of samples with 98.9% accuracy on clean data and 97.9% accuracy under shift. The false-route rate (percentage of misclassified samples among processed ones) increases from 1.1% on clean data to 2.1% under shift. Student1 at this threshold processes 90.2% of samples with 97.6% clean accuracy and 94.6% shifted accuracy, with false-route rates of 2.4% and 5.4% respectively.
In conservative mode (), the Teacher maintains 99.5% accuracy on clean data and 99.0% accuracy under domain shift while processing 89.4% of samples. The false-route rate remains below 1% (0.5% clean, 1.0% shifted), making this operating point particularly suitable for clinical applications where sorting accuracy is paramount. Student1 achieves 98.8% accuracy on clean data and 97.4% under shift at 83.3% coverage, with false-route rates of 1.2% and 2.6% respectively. These results demonstrate that both models can operate in a mode where ambiguous cells are rejected, ensuring that over 98% of processed cells are correctly classified even under domain shift conditions.
Figure S3 synthesizes these findings in a radar chart visualization across four key dimensions: Baseline Accuracy (performance at with 100% coverage), Stability Score (100% minus degradation percentage, quantifying robustness to domain shift), Conservative Accuracy (performance at in high-confidence mode), and Efficiency Score (coverage maintained at ). This multi-dimensional view clearly illustrates that while the Student model achieves remarkable parameter efficiency with a 5000-fold compression, the Teacher maintains superior robustness to domain shift and higher throughput in conservative mode.
From a hardware implementation perspective, the confidence threshold mechanism can be realized with minimal overhead. After the FPGA computes the softmax probabilities, a simple comparator operation checks whether the maximum probability exceeds the threshold . This comparison adds negligible latency (less than 0.1 microseconds) and requires minimal additional logic resources, ensuring that uncertainty-aware rejection does not compromise the sub-25 microsecond total latency budget reported in our main results. The threshold value can be configured as a runtime parameter, allowing system operators to adjust the throughput-safety trade-off based on specific application requirements.
These results establish a quantitative framework for deploying our models in clinical settings where both throughput and safety must be carefully balanced. Research applications prioritizing maximum sample throughput might operate at (over 90% coverage with approximately 98% accuracy), while clinical diagnostic applications demanding highest accuracy would operate at (over 83% coverage with over 98% accuracy). The explicit quantification of false-route rates under domain shift provides critical information for regulatory compliance and risk assessment in clinical deployment scenarios.